THE GROUND STATE FOUR DIMENSIONAL MORPHED POTENTIALS OF HBr AND HI DIMERS

BLAKE A. McELMURRY, ZHONGCHENG WANG, \underline{ROBERT R. LUCCHESE}, and JOHN W. BE-VAN, Chemistry Department, Texas A&M University, College Station, TX 77843-3255; SERGEY P. BELOV, Institute of Applied Physics of Russia Academy of Sciences, Uljanova Str. 46, Nizhny Novgorod 603950, Russia.

The assigned rotational-tunneling transitions including quadrupole substructure of H^{79}Br-H^{81}Br are combined with previously recorded rovibrational infrared and microwave data, to provide a basis for optimizing a ground state four-dimensional morphed potential with quantifiable uncertainties. The determined potential minimum has a value V = -644.0 cm^{−1}at the geometry R = 4.04 Å, θ_{1} = 16.9^{∘}, θ_{2} = 108.3^{∘}, and ϕ = 180.0^{∘}. The peak of the minimum energy path between the two equivalent minima occurs at V = -571.2 cm^{−1}, leading to a barrier height of 72.8 cm^{−1}. A corresponding investigation for the HI dimer will also be presented.